OpenStudy (anonymous):

C=5/9 (F - 32) Part 1: Solve for F Part 2:what is 21 degrees Celsius in ferinhight

OpenStudy (anonymous):

Nevermind that up there^^^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have part 2 done...i just need part 1

OpenStudy (texaschic101):

C = 5/9(F - 32) C = 5/9F - 160/9 (add 160/9 to both sides) C + 160/9 = 5/9F (multiply the equation by 9 to get rid of the fractions) 9C + 160 = 5F (now divide by 5) 9/5C + 32 = F I think

OpenStudy (anonymous):

Anyway to solve for "F" you have to get F to the other side of the equation:

OpenStudy (anonymous):

Can you take a Printscreen of the equation or something? I just want to know the exact equation the way it was.

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

Well the answer is: F = (C x 9/5) + 32

OpenStudy (anonymous):

That's the answer to Part 1.

OpenStudy (anonymous):

ok...thk you i hv #2

OpenStudy (anonymous):

Since you have the formula to find Fahrenheit use it and plug in 21 for C. So: F = (21 x 9/5) + 32 According to PEMDAS do the Parenthesis first: F = 37.8 + 32 Add: 69.8 So 21C = 69.8F

OpenStudy (anonymous):

thk you guys soo very much...you all were so much help and I really apreciate it all...thank you